OpenStudy (anonymous):

Find prime Factorization of 156. Find prime factorization for 306 please show steps

OpenStudy (anonymous):

156 / 2 = 78 78/2 = 39 39/3 = 13 prime factors are 2*2*3*13

OpenStudy (anonymous):

Divide by 2 till you can't anymore, then 3 and further primes until you reach a prime.

OpenStudy (anonymous):

Yeah thanks my algebra book, gives me a really complicated way of doing it

OpenStudy (anonymous):

lets say that 13 was a number like 49 would it be 2*2*3*7*7?

OpenStudy (anonymous):

"13 was a number like 49" ??

OpenStudy (anonymous):

yeah cause in my algebra book it has a different problem and it has the number 49 were the number 13 is in this problem. well this one 882=2*441 =2*3*147 =2*3*3*49 =2*3*3*7*7 =2*3^2*7^2 this is the answer

OpenStudy (anonymous):

882/2, 441/3,147/3,49/7, 7 So u first go with a 2, then a 3 twice, then 7 and end with a 7 is correct. I'm not sure what you mean, you need to check all the possibilities ie if 7 doesn't work, go to the next prime, 11, then 13 etc until u can't do any more.